Given the following information on U.S. Treasury instruments:

1-year note yield = 1.83%                         5-year note yield = 2.79%

2-year note yield = 1.99%                         6-year note yield = 3.04%

3-year note yield = 2.21%                         7-year note yield = 3.58%

4-year note yield = 2.42%                         8-year note yield = 4.18%

And non-changing premiums of 0, .11%, .22%, .39%, .52%, .64%, .75%, .88%, & .98%

Calculate the expected market yield on a 4-year note purchased at the beginning of year 2.
Answer #1

expected market gield on a 4 year note purchased at the begining of year 2: → = 2.42 +0.52 = 2.94-to]